# Emergence of Abstractions: Concept Encoding and Decoding Mechanism for In-Context Learning in Transformers

### v2

@@ Abstract
Humans distill complex experiences into fundamental abstractions that enable rapid learning and adaptation. Similarly, autoregressive transformers exhibit adaptive learning through in-context learning (ICL), which begs the question of how. In this paper, we propose concept encoding-decoding mechanism to explain ICL by studying how transformers form and use internal abstractions in their representations. On synthetic ICL tasks, we analyze the training dynamics of a small transformer and report the coupled emergence of concept encoding and decoding. As the model learns to encode different latent concepts (e.g., “Finding the first noun in a sentence.”) into distinct, separable representations, it concureently builds conditional decoding algorithms and improve its ICL performance. We validate the existence of this mechanism across pretrained models of varying scales (Gemma-2 2B/9B/27B, Llama-3.1 8B/70B). Further, through mechanistic interventions and controlled finetuning, we demonstrate that the quality of concept encoding is causally related and predictive of ICL performance. Our empirical insights shed light into better understanding the success and failure modes of large language models via their representations.
@@ 1 Introduction
Figure 1: An overview of our work. We propose the concept encoding-decoding mechanism to explain why and how task vectors emerge in pretrained LLMs. We demonstrate that transformers concurrently learn to map latent concepts into separable representations and develop context-specific decoding algorithms. We validate the generality of this finding across model families and scales, and show that the quality of concept encoding-decoding can predict ICL task performance.
@@
Hendel et al. (2023); Merullo et al. (2023); Todd et al. (2023) introduce a mechanistic perspective on how pretrained LLMs represent the latent concepts underlying the ICL task as vectors in their representations. They empirically demonstrate that these task-specific vectors can trigger the desired ICL behavior in many cases, with the effectiveness varying across tasks. Although an impactful first observation, there still remains unanswered questions of why these task vectors exist in the first place and why the effectiveness varies by task. This necessitates a deeper mechanistic understanding of the internal abstraction behavior of LLMs, which could explain the findings of task-specific vectors and the workings of ICL.
In our work, we propose the concept encoding-decoding mechanism as the origin of transformers’ internal abstraction behavior. To study the emergence of abstractions during pretraining, we train a small transformer on a mixture of sparse linear regression tasks. We find that concept encoding emerges as the model learns to map different latent concepts into distinct, separable representation spaces. This geometric structuring of the representation space is coupled with the development of concept-specific ICL algorithms – namely, concept decoding. Through causal analysis, we demonstrate that the model associates different algorithms to different learned concepts in the representation space and that ICL happens through the two-step process. Importantly, we see that the emergence of the two-stage process coincides with each other, suggesting a mutual dependence between the two.
We demonstrate the validity of the concept encoding-decoding mechanism across different pretrained model families and scales (Llama-3.1-8B/70B and Gemma-2 2B/9B/27B) on more natural ICL tasks, such as part-of-speech tagging and bitwise arithmetic. We show that large language models (LLMs) trained on diverse data also exhibit concept encoding-decoding behavior. With more in-context examples, LLMs map the inputs to increasingly separable representation spaces, clustered by the latent concepts. Moreover, leveraging insights from the synthetic experiments, we demonstrate that the decodability of the concepts from representations is predictive of downstream ICL performance. We establish a causal relationship between the concept decodability and ICL performance through mechanistic intervention and controlled finetuning experiments.
@@
    We first study the emergence of task vectors by training a small transformer on a synthetic ICL task (§3.3) and propose concept encoding-decoding mechanism to explain the emergent behavior for learning to solve ICL tasks. We observe that earlier layers of the model learn to encode the latent concept, whereas the latter layers conditions the algorithm on the inferred concept. Interestingly, the emergence of the two-stage process is coupled, implying a mutual dependence.
@@
    We introduce Concept Decodability (CD) as a geometric measure of internal abstraction formation for latent concepts and demonstrate that CD effectively predicts downstream ICL performance in pretrained LLMs (§4.2). We demonstrate our framework’s generality across tasks, model families, and scales (Llama 3.1 8B/70B, Gemma 2B/9B/27B).
@@
    We establish the causal relationship between CD and ICL performance in pretrained LLMs through mechanistic intervention (§4.1) and controlled finetuning (§4.3).
@@ 2 Related Work
Astounded by LLMs’ ability to perform ICL, many have proposed theories to understand the mechanisms of ICL. Some (Dai et al., 2023; von Oswald et al., 2023; Ahn et al., 2024; Akyürek et al., 2024) have proposed that LLMs, with linear attention (Katharopoulos et al., 2020), can implement stochastic gradient descent to perform ICL. Other works (Xie et al., 2021; Wang et al., 2024; Ye et al., 2024) have presented a Bayesian framework to theoretically explain the workings of ICL. This view implies that the model implements a two-stage algorithm to estimate the posterior $P(z|\mathcal{D})$ and the likelihood $P(y_{\ast}|x_{\ast},\mathcal{D})$. In this work, we adopt this framework and demonstrate how the model implements it through its intermediate representations. More specifically, we study the emergence of the concept encoding-decoding mechanism.
Task Vectors.
Todd et al. (2023) and Hendel et al. (2023) identify task-specific vectors in LLMs that can induce desired ICL task behavior (e.g., object-color mapping). Building on these work, our study explores the origin of such vectors and describes how they naturally emerge within transformer through the concept encoding-decoding mechanism. Moreover, we adopt similar mechanistic intervention tools from their work in Section 4.2. However, beyond simply showing the existence of task-specific vectors, we leverage the intervention studies to demonstrate a causal relationship between concept encoding and the concept-specific decoding algorithms.
Latent Concepts in Language Model Representations.
Several studies (Dalvi et al., 2022; Merullo et al., 2023) have examined how language models encode concepts in their representations. In LLMs, notions, such as truthfulness (Marks & Tegmark, 2023), time, and space(Gurnee & Tegmark, 2024), have been shown to be linearly separable representations. Sparse autoencoders (Bricken et al., 2023; Cunningham et al., 2023) have revealed highly interpretable features emerge with scale in LLMs. Beyond the identification of these concepts, our work aims to answer how such concepts emerge in the representation of LLMs and how they can predictive downstream ICL task performance.
@@
To study the causal relationship between the quality of concept encoding-decoding and downstream ICL performance, we adopt causal mediation analysis techniques from Geiger et al. (2020); Vig et al. (2020); Todd et al. (2023); Heimersheim & Nanda (2024); Merullo et al. (2024). We specifically use the method of activation patching, where we replace the activations of an immediate layer from a sample with another. This technique allows us to demonstrate that transformers implement different algorithms conditioned on the inferred concepts.
@@ 3.1 Notation and Background
We focus on ICL problems, where the goal is to predict $y_{\ast}$ from a query $x_{\ast}$, given some in-context examples $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$. Each problem shares a latent concept $\bm{z}$ that links inputs $\bm{x}$ to outputs $\bm{y}$. For instance, in an ICL task where latent concept is object-color mapping, we provide demonstrations like (apple, red), (banana, yellow), and (grape, purple), and then ask for what comes after (lemon, ?). We employ this parameterization to accommodate latent concepts varying in complexity, from simple function regression problems (Garg et al., 2022; von Oswald et al., 2023; Li et al., 2023) to POS tagging (Blevins et al., 2022; Banko & Moore, 2004) and arithmetic (He et al., 2024).
@@ 3.2 Theoretical Framework
Although theoretically compelling, it was not until recently that Hendel et al. (2023); Todd et al. (2023); Merullo et al. (2023) showed empirical evidence of models encoding the latent concepts in the intermediate representations. They illustrate that concept-specific vectors are then decoded and trigger the desired ICL task behavior. With simple encoder-decoder analogy, these findings suggest that the two-stage behavior of ICL, as described in Eq. 1, is mediated by the encoding and decoding of latent variables within the representation space. Building on this idea, we begin our investigation with the following questions:
@@
    How does the concept encoding-decoding behavior emerge in the model over training and how do they interplay?
@@
    How is the model’s ability to accurately infer the latent concepts related to downstream ICL performance?
@@ 3.3 Motivation: Synthetic Experiments
Figure 2: Coupled emergence of concept encoding and conditional decoding algorithms in mixture of sparse linear regression. The loss curve on the left-hand side shows different convergence dynamics per basis and show three phases of descent, which we mark with (a), (b), and (c). On the right-hand side, we plot the geometric changes in the representations and how they separate by basis at these marked points. These points coincide with the algorithmic switching behavior.
We train a small transformer on a synthetic ICL task and demonstrate that concept encoding and decoding emerges simultaneously during training. Through causal analysis, we show that, as the models “discovers” a latent concept by building a distinct representation space from the others, it associates the concept with a uniquely corresponding decoding algorithm. Finally, we propose the concept encoding-decoding mechanism that encompasses these findings and serve as the core theory throughout the remainder of our study.
@@
We compose our task as a mixture of sparse linear regression tasks. We follow the conventional linear regression setup from Garg et al. (2022); von Oswald et al. (2023) and construct the input-output pair $(x_{i},y_{i})$ by sampling $x_{i}\sim\mathcal{N}(0,\bm{I_{D}})$ and $y_{i}=W^{T}x_{i}+\epsilon_{i}$, where $W$ is randomly generated from a standard normal distribution, $\mathcal{N}(0,\bm{I_{D}})$, and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. We, however, add sparsity constraints to $W$ with the sparsity pattern represented by the basis $B_{k}$. Each $B_{k}$ has a rank of $r$. In other words, the basis chooses the dimensions of $W$ to turn on and off. The basis is sampled uniformly from =$\mathcal{B}=\{B_{1},B_{2},B_{3},B_{4}\}$ and each basis is non-overlapping and orthogonal to each other. By default, we set $D=16$ and $r=4$. By adding this layer of latent concept of $\bm{B}$, we can explicitly control and interpret the latent concepts, and analyze their representations.
@@
We train a 12-layer GPT-2 architecture transformer (Radford et al., 2019) with an embedding dimension of 128 and 8 attention heads. We train the model to minimize mean squared error (MSE) loss over the sequence length of 20. We run 5 different random seeds for training and report observations that generalize across the runs. We detail the experimental setup further in Appendix C.2.
@@
The error bounds of our task depend on whether the model learns to infer the underlying bases. If the model learns to infer the bases, then the model can achieve $r$-dimensional regression, where the MSE approaches 0 with $r$ in-context examples. If not, the model, in the worst case, can perform $D$-dimensional regression with $r$-sparsity, which has a longer tailed error curve that approaches 0 between $r$ and $D$ in-context examples. With these insights, we can better analyze which latent basis the model has learned and the associated algorithm. Note that we define “algorithm” as a class of statistical methods for linear regression, as detailed in Appendix C.1.
@@
Observation 1: Different Loss Dynamics Per Basis. We interestingly observe that each basis, despite having identical task complexities, exhibits different loss descents during training. Figure 2 shows the test MSE averaged over the sequence over training. $B_{1}$ displays a distinct loss descent dynamic, undergoing an abrupt drop at epoch 10. In contrast, the other three bases, $B_{2}$, $B_{3}$, and $B_{4}$, exhibit correlated loss descent dynamics, with two smaller descents at 10 and 40 epochs. This suggests that the model learns to infer $B_{1}$ differently and applies selective algorithms.
@@
Observation 2: Emergence of Separable Representations and Coupled Algorithmic Phase Transitions. We also analyze the geometry of the intermediate representations at layer 5 to question how the model may be encoding the latent bases. Surprisingly, at the three points of descent (a, b, c) marked in Figure 2, the model gradually builds separate representations for the different bases as shown in the UMAP visualizations. At point (a), the three bases are clustered together and the model’s algorithm resembles a 16-dimensional weighted LASSO regression. As $B_{1}$ separates out at point (b), the model starts to leverage the inferred basis to switch to a 4-dimensional regression. At point (c), when all four classes are separable, the model converges to the optimal 4-dimensional regression. This observation suggests that model encode latent concepts into separated representations to conditionally apply decoding algorithms.
Causal Relation between Concept Encoding and Performance.
@@
We conduct perturbation analysis to validate that the model conditionally applies decoding algorithms based on the separated representations. Given an input of a source basis, we patch the activations of layer 5 – representations of the residual stream of the transformer layer – with the mean activations of a target basis and analyze whether it will improve or degrade performance. When the source is equal to the target (self-perturbation), the patching should help the model identify the basis and improve performance. Otherwise, it should hinder correct basis inference and therefore performance. We perform this analysis at points (b) and (c) from Figure 2, when the latent concept representations are semi and fully separable.
@@
In Figure 3, we present the perturbation analysis at point (b) on the left. In this case, $B_{2,3,4}$ forms one cluster and $B_{1}$ another. We observe that all the self-perturbations along the diagonal and intracluster ($B_{2,3,4}$) slightly decrease the loss or show no effect. However, when we apply perturbations across different clusters, the loss spikes, indicating that we trigger different decoding algorithms unsuitable for the input sequence. This analysis shows that, because the model was only able to encode two different latent concepts in the intermediate representations, it only learns two classes of algorithms, one for $B_{1}$ and another for $B_{2,3,4}$.
@@
On the right of Figure 3, we conduct the same perturbation study at convergence, when the model learns to encode all of the latent concepts as distinct representations. Surprisingly, we observe that the model undergoes an algorithmic phase transition of implementing concept-conditioned algorithms. Not only does all the self-perturbation along the diagonal improve performance more noticeably, but also any perturbation to a different basis results in significantly higher losses.
@@
These results altogether draw the picture that a transformer, when trained to perform ICL, gradually learns to encode the latent concepts to separable representation spaces and learns to conditionally apply decoding algorithms simultaneously. These observations suggest that concept encoding and decoding are mutually dependent. It is possible that they are mutually reinforcing as well, but this needs further investigation.
Generalizability to More Complex Synthetic Tasks
@@
We also question whether the observed co-emergence of concept-encoding and decoding can robustly hold in more complex settings. To this end, we perform ablation studies in the sparse linear regression setup, where we increase the number of basis functions and introduce non-orthogonal bases with overlaps. We describe the specific details in Appendix D. Although our core observation, the coupled emergence of the concept encoding-decoding mechanism, is also evident in these complex tasks, an intriguing finding emerges: bases that share overlap and are correlated, even at convergence, are not fully separated in the representation space and share the same loss over training. This suggests that in natural text, where concepts are often correlated and exhibit semantic overlaps, the representation space within models may similarly fail to fully disentangle these concepts, potentially leading to shared representational subspaces that could impact the model’s ability to generalize or distinguish between subtly different but related tasks.
3.4 Concept Encoding-Decoding Mechanism

### v3

# Emergence and Effectiveness of Task Vectors in In-Context Learning: An Encoder Decoder Perspective

@@ Abstract
Autoregressive transformers exhibit adaptive learning through in-context learning (ICL), which begs the question of how. Prior works have shown that transformers represent the ICL tasks as vectors in their representations. In this paper, we leverage the encoding-decoding framework to study how transformers form task vectors during pretraining and how their task encoding quality predicts ICL task performance. On synthetic ICL tasks, we analyze the training dynamics of a small transformer and report the coupled emergence of task encoding and decoding. As the model learns to encode different latent tasks (e.g., “Finding the first noun in a sentence.”) into distinct, separable representations, it concurrently builds conditional decoding algorithms and improves its ICL performance. We validate this phenomenon across pretrained models of varying scales (Gemma-2 2B/9B/27B, Llama-3.1 8B/70B) and over the course of pretraining in OLMo-7B. Further, we demonstrate that the quality of task encoding inferred from representations predicts ICL performance, and that, surprisingly, finetuning the earlier layers can improve the task encoding and performance more than finetuning the latter layers. Our empirical insights shed light into better understanding the success and failure modes of large language models via their representations.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: An overview of our work. We study the task encoding-decoding to explain why and how task vectors emerge in pretrained LLMs. We demonstrate that transformers concurrently learn to map latent concepts into separable representations and develop task-specific decoding algorithms. We validate the generality of this finding across model families and scales, and show that the quality of task encoding-decoding can predict ICL task performance.
@@
Hendel et al. (2023); Merullo et al. (2023); Todd et al. (2023) introduce a mechanistic perspective on how pretrained LLMs represent the latent concepts underlying the ICL task as vectors in their representations. They empirically demonstrate that these task-specific vectors can trigger the desired ICL behavior in many cases, with the effectiveness varying across tasks. Although an impactful first observation, there still remains unanswered questions of why these task vectors exist in the first place and why the effectiveness varies by task. This necessitates a deeper mechanistic understanding of this internal abstraction behavior of LLMs.
In our work, we leverage the encoding-decoding framework to investigate the origin and formation of task vectors in transformers. To study their emergence during pretraining, we train a small transformer on a mixture of sparse linear regression tasks. We find that task encoding emerges as the model learns to map different latent tasks into distinct, separable representation spaces. This geometric structuring of the representation space is coupled with the development of task-specific ICL algorithms – namely, task decoding. Through causal analysis, we demonstrate that the model associates different algorithms to different learned concepts in the representation space and that ICL happens through the two-step process. Importantly, we see that the emergence of the two-stage process coincides with each other, suggesting a mutual dependence between the two.
Inspired by these findings, we investigate the task encoding-decoding phenomenon across different pretrained model families and scales (Llama-3.1-8B/70B and Gemma-2 2B/9B/27B) on more natural ICL tasks, such as part-of-speech tagging and bitwise arithmetic. We introduce task decodability, a geometric measure that quantifies how effectively a model can decode the task from its intermediate representations. We demonstrate that task decodability is causally related to ICL performance through ablation studies on ICL examples, finetuning, and prompting. Overall, our study reveals that the encoder-decoder frameworks provide valuable insights into the emergence of task vectors and why their effectiveness varies across tasks.
@@
    We first study the emergence of task vectors during training under the encoder-decoder framework. By training a small transformer on synthetic ICL tasks (§3.3), we observe the coupled occurrence of task encoding and task decoding, ultimately forming task vectors.
@@
    We introduce Task Decodability (TD) as a geometric measure to quantify how well the model can infer the latent tasks from its representations and demonstrate that TD effectively predicts downstream ICL performance in pretrained LLMs (§4.2). We demonstrate our framework’s generality across tasks, model families, and scales (Llama 3.1 8B/70B, Gemma 2B/9B/27B) and further study its evolution throughout pretraining using different checkpoints of OLMo-7B.
@@
    We establish the causal relationship between TD and ICL performance in pretrained LLMs through mechanistic intervention (§4.1) and controlled finetuning (§4.4). Contrary to common convention, we show that finetuning the earlier layers improves ICL performance more than finetuning the later layers (Wu et al., 2024; Kumar et al., 2022).
@@ 2 Related Work
Astounded by LLMs’ ability to perform ICL, many have proposed theories to understand the mechanisms of ICL. Some works (Dai et al., 2023; von Oswald et al., 2023; Ahn et al., 2024; Akyürek et al., 2024) have proposed that LLMs, with linear attention (Katharopoulos et al., 2020), can implement stochastic gradient descent to perform ICL. Other works (Xie et al., 2021; Wang et al., 2024; Ye et al., 2024) have presented a Bayesian framework to theoretically explain the workings of ICL. This view implies that the model implements a two-stage algorithm to estimate the posterior $P(z|\mathcal{D})$ and the likelihood $P(y_{\ast}|x_{\ast},\mathcal{D})$. In this work, we adopt this framework and demonstrate how the model implements it through its intermediate representations. More specifically, we study the emergence of task encoding and decoding.
Task Vectors and Latent Concepts in LLM Representations.
Recent work by Todd et al. (2023) and Hendel et al. (2023) identifies task-specific vectors in LLMs that can induce desired in-context learning behaviors (e.g., object-color mapping). Building on this foundation, our study examines when and how task-specific representations emerge and how their quality (TD score) can be measured and used to predict downstream ICL performance. Moreover, Park et al. (2024) study the different algorithmic stages for ICL under homogeneous Markov chain settings. In contrast, our work expands the study to heterogeneous tasks, a regime more reflective of LLM pretraining, and proposes a more general encoding-decoding framework to understand the mechanics of ICL.
Beyond the scope of task-specific vectors, several studies have explored how language models encode a wide range of latent concepts (Dalvi et al., 2022; Merullo et al., 2023), including truthfulness (Marks & Tegmark, 2023), time, and space (Gurnee & Tegmark, 2024). These investigations reveal that such notions can be linearly separable in the hidden representations, and that model scaling often yields more interpretable features (Bricken et al., 2023; Cunningham et al., 2023)
@@
To study the causal relationship between the quality of task encoding-decoding and downstream ICL performance, we adopt causal mediation analysis techniques from Geiger et al. (2020); Vig et al. (2020); Todd et al. (2023); Heimersheim & Nanda (2024); Merullo et al. (2024). We specifically use the method of activation patching, where we replace the activations of an intermediate layer from a sample with another.
@@ 3.1 Notation and Background
We focus on ICL problems, where the goal is to predict $y_{\ast}$ from a query $x_{\ast}$, given some in-context examples $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$. Each problem shares a latent task $\bm{z}$ that links inputs $\bm{x}$ to outputs $\bm{y}$. For instance, in an ICL task where the latent task is object-color mapping, we provide demonstrations like (apple, red), (banana, yellow), and (grape, purple), and then ask for what comes after (lemon, ?). We employ this parameterization to accommodate latent tasks varying in complexity, from simple function regression problems (Garg et al., 2022; von Oswald et al., 2023; Li et al., 2023) to Part-of-Speech (POS) tagging (Blevins et al., 2022; Banko & Moore, 2004) and bitwise arithmetic (He et al., 2024).
@@ 3.2 Theoretical Framework
Although theoretically compelling, it was not until recently that Hendel et al. (2023); Todd et al. (2023); Merullo et al. (2023) showed empirical evidence of models encoding the latent concepts in the intermediate representations. They illustrate that task-specific vectors are then decoded and trigger the desired ICL task behavior. With a simple encoder-decoder analogy, these findings suggest that the two-stage behavior of ICL, as described in Equation 1, is mediated by the encoding and decoding of latent variables within the representation space. Building on this idea, we begin our investigation with the following questions:
@@
    How do task vectors emerge during training, and what drives their varying effectiveness across tasks?
@@
    How is the model’s ability to accurately infer the latent tasks related to downstream ICL performance?
@@ 3.3 Motivation: Synthetic Experiments
Figure 2: Coupled emergence of task encoding and conditional decoding algorithms in mixture of sparse linear regression. The loss curve on the left-hand side shows different convergence dynamics per basis and show three phases of descent, which we mark with (a), (b), and (c). The test MSE is the mean squared error computed over a sequence of 20 in-context examples. On the right-hand side, we plot the geometric changes in the representations and how they separate by basis at these marked points. These points coincide with the algorithmic switching behavior. For the UMAP visualizations, we randomly draw 100 samples for each basis. Next, we collected $y$-token representations of 5th layer at 20th demonstrations and plotted the UMAP with the parameters of n neighbors$=$15 and min dist$=$0.1.
As a motivating experiment, we study the formation of task vectors during training dynamics of a small autoregressive transformer on synthetic ICL tasks. We observe that, as the model “localizes” the latent task by building a distinct representation from the others, it associates it with a uniquely corresponding decoding algorithm. Building on this observation, we outline the task encoding and decoding framework to explain the formation and mechanism of task vectors.
@@
We compose our task as a mixture of sparse linear regression. We follow the conventional linear regression setup from Garg et al. (2022); von Oswald et al. (2023) and construct the input-output pair $(x_{i},y_{i})$ by sampling $x_{i}\sim\mathcal{N}(0,\bm{I_{D}})$ and $y_{i}=W^{T}x_{i}+\epsilon_{i}$, where $W$ is randomly generated from a standard normal distribution, $\mathcal{N}(0,\bm{I_{D}})$, and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$. We, however, add sparsity constraints to $W$ with the sparsity pattern represented by the basis $B_{k}$. Each $B_{k}$ has a rank of $r$. In other words, the basis chooses the dimensions of $W$ to turn on and off. The basis is sampled uniformly from =$\mathcal{B}=\{B_{1},B_{2},B_{3},B_{4}\}$ and each basis is non-overlapping and orthogonal to each other. By default, we set $D=16$ and $r=4$. By adding this layer of latent concept of $\bm{B}$, we can explicitly control and interpret the latent concepts, and analyze their representations.
@@
We train a 12-layer GPT-2 architecture transformer (Radford et al., 2019) with an embedding dimension of 128 and 8 attention heads. We train the model to minimize mean squared error (MSE) loss over the sequence length of 20. We run 5 different random seeds for training and report observations that generalize across the runs. We detail the experimental setup in Appendix D.2.
@@
The error bounds of regression depend on whether the model learns to infer the underlying bases. If the model can infer the bases, then the model can theoretically achieve $r$-dimensional regression, where the MSE approaches 0 with $r$ in-context examples. If not, the model, in the worst case, can perform $D$-dimensional regression with $r$-sparsity, which has a longer tailed error curve that approaches 0 with $D$ in-context examples. With these insights, we can better analyze which latent basis the model has learned and the associated algorithm from its error curve. Note that we define “algorithm” as a class of statistical methods for linear regression, as detailed in Appendix D.1.
@@
Observation 1: Different Loss Dynamics per Basis. We interestingly observe that each basis, despite having identical task complexities, exhibits different loss descents during training. Figure 2 shows the test MSE averaged over the sequence over training. $B_{1}$ displays a distinct loss descent dynamic, undergoing an abrupt drop at epoch 10. In contrast, the other three bases, $B_{2}$, $B_{3}$, and $B_{4}$, exhibit correlated loss descent dynamics, with two smaller descents at 10 and 40 epochs. This suggests that the model learns to infer $B_{1}$ differently and applies a selective decoding algorithm.
@@
Observation 2: Emergence of Separable Representations and Coupled Algorithmic Phase Transitions. We also analyze the geometry of the intermediate representations at layer 5 to question how the model may be encoding the latent bases. Surprisingly, at the three points of descent (a, b, c) marked in Figure 2, the model gradually builds separate representations for the different bases as shown in the UMAP visualizations. At point (a), the three bases are clustered together and the model’s algorithm resembles a 16-dimensional weighted LASSO regression. As $B_{1}$ separates out at point (b), the model starts to leverage the inferred basis to switch to a 4-dimensional regression. At point (c), when all four classes are separable, the model converges to the optimal 4-dimensional regression. This observation suggests that model encodes the tasks into separated representations to conditionally apply decoding algorithms.
Causal Relation between Task Encoding and Performance.
@@
We conduct perturbation analysis to validate that the model conditionally applies decoding algorithms based on the separated representations. Given an input of a source basis, we patch the activations of layer 5 – representations of the residual stream of the transformer layer – with the mean activations of a target basis and analyze whether it will improve or degrade performance. We specify layer 5 for this activation patching analysis because the separation of representations by concept is only clearly observed from that layer and afterwards based on the UMAP visualizations in Figure 11. We formally describe the activation patching procedure in Appendix C. When the source is equal to the target (self-perturbation), the patching should help the model identify the basis and improve performance. Otherwise, it should hinder correct basis inference and therefore degrade performance. We perform this analysis at different steps of training – (b) and (c) from Figure 2, when the latent task representations are semi and fully separable.
@@
In Figure 7 of Appendix D.3, we present the perturbation analysis at point (b) on the left. In this case, $B_{2,3,4}$ forms one cluster and $B_{1}$ another. We observe that all the self-perturbations along the diagonal and intracluster ($B_{2,3,4}$) slightly decrease the loss or show no effect. However, when we apply perturbations across different clusters, the loss spikes, indicating that we trigger different decoding algorithms unsuitable for the input sequence. This analysis shows that, because the model was only able to encode two different latent concepts in the intermediate representations, it only learns two classes of algorithms, one for $B_{1}$ and another for $B_{2,3,4}$.
@@
On the right of Figure 7, we conduct the same study at convergence, when the model learns to encode all of the latent concepts as distinct representations. Surprisingly, we observe that the model undergoes an algorithmic phase transition of implementing individual task-specific algorithms. Not only does all the self-perturbation along the diagonal improve performance more noticeably, but also any perturbation to a different basis results in significantly higher losses.
@@
These results altogether draw the picture that a transformer, when trained to perform ICL, gradually learns to encode the latent tasks into separable representation spaces and learns to conditionally apply decoding algorithms simultaneously. These observations suggest that task encoding and decoding are mutually dependent.
Generalizability to Increasing Dimensions and Basis Overlap
@@
We investigate this coupled emergence of task encoding and decoding under more complex settings. To this end, we increase the number of tasks and introduce non-orthogonal bases with overlaps (i.e., some bases are correlated). We describe the specific details in Appendix E. First, we, once again, observe the coupled emergence of task encoding and decoding in these more complex tasks in Figure 13 of Appendix E. Second, as shown in Figure 15 of Appendix E, a more intriguing pattern emerges: bases that share overlap and are correlated, even at convergence, are not fully separated in the representation space and share the same loss over training. This suggests that, in natural text, where tasks are often correlated and exhibit semantic overlaps, the model may similarly fail to disentangle their representations and distinguish these tasks.
Generalizability to Mixture of Various Regression Tasks
To further validate the robustness of the task encoding-decoding framework, we trained a transformer using a mixture of various regression tasks (linear, sparse linear, leaky ReLU, and quadratic regression) from Li et al. (2023) detailed in Appendix F. In Figure 16, the model successfully learns linear, sparse linear, and leaky ReLU regression but fails at quadratic regression. Despite strong ICL performance, intermediate representations across the regression tasks in intermediate layers have high overlap. We hypothesize that task vectors do not naturally emerge here because these tasks could be solved with the same core linear regression algorithm (Li et al., 2023).
To verify the conjecture, we perform attention head pruning experiments detailed in Appendix F and report that the different regression families, excluding quadratic regression, share the same decoding algorithm. We sequentially prune each attention head and measure the resulting change in MSE. Figure 17 shows that in the mixture of regression tasks, there is a consistent sharing of attention heads across linear regression, leaky ReLU, and sparse linear regression tasks, excluding quadratic regression. This contrasts with the head pruning result of the sparse linear regression in Figure 18, where each basis seemingly operates with a different set of attention heads and algorithm. This suggests structural similarity in the algorithms in this new mixture setting and provides a direct evidence that they are structurally indistinguishable within the model, supporting our hypothesis.
3.4 Task Encoding-Decoding